# Parity-Aware Byte-Pair Encoding: Improving Cross-lingual Fairness in Tokenization

### v1

@@ Abstract
Tokenization is the first—and often least scrutinized—step of most NLP pipelines. Standard algorithms for learning tokenizers rely on frequency-based objectives, which favor languages dominant in the training data and consequently leave lower-resource languages with tokenizations that are disproportionately longer, morphologically implausible, or even riddled with <UNK> placeholders. This phenomenon ultimately amplifies computational and financial inequalities between users from different language backgrounds. To remedy this, we introduce Parity-aware Byte Pair Encoding (BPE), a variant of the widely-used BPE algorithm. At every merge step, Parity-aware BPE maximizes the compression gain of the currently worst-compressed language, trading a small amount of global compression for cross-lingual parity. We find empirically that Parity-aware BPE leads to more equitable token counts across languages, with negligible impact on global compression rate and no substantial effect on language-model performance in downstream tasks. [^1] Code available here.
@@ 1 Introduction
At a time of rapid innovation and constant change in natural language processing (NLP), tokenization continues to be a foundational and comparatively stable component of NLP pipelines. Tokenization is the transformation of raw sequences of bytes [^2] Early work considered characters the “base unit” of strings, but bytes have become popular because their fixed 256-symbol vocabulary can encode any character from any encoding, eliminating out-of-vocabulary issues. into sequences of byte-spans, i.e., subwords; it enables computational efficiency and provides essential inductive biases by defining meaningful textual units. This design choice can have a major impact on various aspects of model performance (Bostrom and Durrett, 2020; Ali et al., 2024; Goldman et al., 2024).
@@
In the effort to mitigate these inequities, we introduce Parity-aware BPE. The classic version of BPE learns its vocabulary by repeatedly selecting the subword pair with the highest corpus-level co-occurrence count; it adds the concatenation of this pair to the vocabulary and replaces all pair co-occurrences with the new symbol. [^3] This iterative merging process can be viewed as a form of data compression, where frequent subword sequences are replaced with shorter representations. Parity-aware BPE is a simple variant of this algorithm, retaining the iterative framework but redefining the merge selection rule: at each step, it computes co-occurrence statistics separately for each language and then uses statistics from the language with the current worst compression rate for selecting the next merge. In other words, instead of greedily maximizing a global objective, Parity-aware BPE performs a “fair-max” update that targets the worst-off language. It thus progressively equalizes string compression rates across languages.
@@ 2.2 Text Compression
Many commercial technology services offer APIs that bill per token; a service’s processing speed also scale with the number of tokens in the input and output. Token counts thus dictate both the economics and the latency of these services. For a given byte sequence $\mathbf{b}$, the number of tokens produced by tokenization function $\tau$ is determined by the tokenizer’s compression rate $\mathrm{CR}(\mathbf{b};\tau)$. $\mathrm{CR}(\ell;\tau)$—the expected compression rate over sequences from language $\ell$—is thus a direct proxy for the average per-language cost (and expected latency) of using one of these services. Variance in $\mathrm{CR}(\ell;\tau)$ across languages thus implies different user costs purely as a function of language choice (Petrov et al., 2023), and whether $\mathrm{CR}(\ell;\tau)$ is comparable across languages is therefore one way of measuring tokenizer fairness. Byte Pair Encoding, discussed next, optimizes for compression rate across a corpus without regard for language. Our subsequent adjustment to Classical BPE adds an auxiliary objective of equalizing $\mathrm{CR}(\ell;\tau)$ across languages.
@@ 2.3 Byte Pair Encoding
[…]
$\text{Step 2}:\texttt{ba b}\!\to\!\texttt{bab}\,\,\Longrightarrow\,\,\texttt{ba bab}$
@@ 3.3 Algorithmic Variants
Model developers may want to include and tokenize data for which parallel data is not available or where this concept does not even apply, such as programming code. Also, they may not want to guarantee full parity, but still give a high weight to global compression. We support these goals with a hybrid learning algorithm that uses the global objective of vanilla BPE (eq.˜5) for the first $K$ merges, then switches to the parity-aware objective (eq.˜6) for another $J$ merges. $K$ and $J$ can be chosen by model developers to assign budgets to global compression and fairness.
@@ 4.2.1 Intrinsic Metrics
  • Tokenizer Fairness Gini adapts the Gini coefficient to the per-language tokenization cost distribution (e.g., tokens per line (document) in a parallel corpus). Values near 0 mean equal cost across languages; values closer to 1 indicate inequality.
@@
For completeness, we also track Type–Token Ratio and Average Token Rank (vocabulary diversity; Limisiewicz et al., 2023) as well as Rényi entropies (distributional concentration; Zouhar et al., 2023a).
@@ 5.1 Intrinsic Evaluation
Table 2: Average LM performance (accuracy %) across 13 multilingual benchmarks (tokenizers trained on the unbalanced 30-lang dataset with $128k$ vocab size). The Random column shows the expected accuracy of a random classifier. Best performance per language is bolded. Benchmark descriptions and details for each language are provided in ˜E and 7, respectively.
@@ 7 Discussion and Conclusion
Overall, the trade-offs required for using parity-aware BPE are minimal. During the learning stage, the algorithm adds only an $\mathcal{O}(\mathcal{L})$ pass per merge for recomputing language-level compression rates on a dev corpus, leaving the asymptotic complexity identical to Classical BPE. Further, we observe empirically that a small, sentence-aligned development set is sufficient to drive the “fair-max” decision. When resource or domain mismatches make full equality undesirable, hybrid and moving-window variants further let practitioners trade off global compression versus strict parity; our empirical results validate that these variants perform well in practice. From the model-developer’s perspective, parity-aware BPE is a drop-in replacement: it requires no architectural changes and minimal changes to the tokenizer pipeline.
@@ Acknowledgements
RS acknowledges support by the Swiss National Science Foundation through the MUTAMUR project (no. 213976). We also gratefully acknowledge the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, a Meta LLM Evaluation Research Grant, and the Swiss AI Initiative through a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a06 on Alps.
@@ B.3 Morphological and Multilingual Fairness Metrics
We adapt the Gini coefficient to measure fairness across languages by treating token costs (fertility values) as a distribution: Let $c_{1}\leq c_{2}\leq\ldots\leq c_{n}$ be the “costs” for languages $\mathcal{L}=\{l_{1},l_{2},\ldots,l_{n}\}$ incurred by a given tokenizer. Here, we quantify cost as the average number of tokens it takes to encode the unit of interest (e.g., a byte, word or line); when using a parallel corpus, this can be cost per line (document), which controls for discrepancies between average character byte length across different scripts. The Gini coefficient is then:
@@
Values range from 0 (perfect equality) to 1 (maximum inequality). This metric condenses multilingual tokenizer fairness into a single number by measuring the degree of inequality in computational costs across languages; lower Gini coefficients indicate more equitable tokenizer compression across languages, while higher values suggest systematic bias toward certain languages.

### v2

@@ Abstract
Tokenization is the first—and often least scrutinized—step of most NLP pipelines. Standard algorithms for learning tokenizers rely on frequency-based objectives, which favor languages dominant in the training data and consequently leave lower-resource languages with tokenizations that are disproportionately longer, morphologically implausible, or even riddled with <UNK> placeholders. This phenomenon ultimately amplifies computational and financial inequalities between users from different language backgrounds. To remedy this, we introduce Parity-aware Byte Pair Encoding (BPE), a variant of the widely-used BPE algorithm. At every merge step, Parity-aware BPE maximizes the compression gain of the currently worst-compressed language, trading a small amount of global compression for cross-lingual parity. We find empirically that Parity-aware BPE leads to more equitable token counts across languages, with negligible impact on global compression rate and no substantial effect on language-model performance in downstream tasks.
[Uncaptioned image] swiss-ai/parity-aware-bpe
@@ 1 Introduction
At a time of rapid innovation and constant change in natural language processing (NLP), tokenization continues to be a foundational and comparatively stable component of NLP pipelines. Tokenization is the transformation of raw sequences of bytes [^1] Early work considered characters the “base unit” of strings, but raw bytes have become popular because their fixed 256-symbol vocabulary can encode any character from any encoding, eliminating out-of-vocabulary issues. into sequences of byte-spans, i.e., subwords; it enables computational efficiency and provides essential inductive biases by defining meaningful textual units. This design choice can have a major impact on various aspects of model performance (Bostrom and Durrett, 2020; Ali et al., 2024; Goldman et al., 2024).
@@
In the effort to mitigate these inequities, we introduce Parity-aware BPE. The classic version of BPE learns its vocabulary by repeatedly selecting the subword pair with the highest corpus-level co-occurrence count; it adds the concatenation of this pair to the vocabulary and replaces all pair co-occurrences with the new symbol. [^2] This iterative merging process can be viewed as a form of data compression, where frequent subword sequences are replaced with shorter representations. Parity-aware BPE is a simple variant of this algorithm, retaining the iterative framework but redefining the merge selection rule: at each step, it computes co-occurrence statistics separately for each language and then uses statistics from the language with the current worst compression rate for selecting the next merge. In other words, instead of greedily maximizing a global objective, Parity-aware BPE performs a “fair-max” update that progressively equalizes string compression rates across languages. Notably, this modification affects only vocabulary learning; the inference procedure remains the same as in Classical BPE.
@@ 2.2 Text Compression
Many commercial technology services offer APIs that bill per token; services’ processing speeds also scale with the number of tokens in the input and output. Token counts thus dictate both the economics and the latency of these services. For a given byte sequence $\mathbf{b}$, the number of tokens produced by tokenization function $\tau$ is determined by the tokenizer’s compression rate $\mathrm{CR}(\mathbf{b};\tau)$. $\mathrm{CR}(\ell;\tau)$—the expected compression rate over sequences from language $\ell$—is thus a direct proxy for the average per-language cost (and expected latency) of using one of these services. Variance in $\mathrm{CR}(\ell;\tau)$ across languages thus implies different user costs purely as a function of language choice (Petrov et al., 2023), and whether $\mathrm{CR}(\ell;\tau)$ is comparable across languages is therefore one way of measuring tokenizer fairness. Byte Pair Encoding, discussed next, optimizes for compression rate across a corpus without regard for language. Our subsequent adjustment to Classical BPE adds an auxiliary objective of equalizing $\mathrm{CR}(\ell;\tau)$ across languages.
@@ 2.3 Byte Pair Encoding
$\mathbf{m}=$ $[(\texttt{b},\texttt{a}),(\texttt{ba},\texttt{b})];\quad\mathbf{b}=\texttt{babab}\phantom{a}$
$\quad\mathbf{v}_{0}=\texttt{b},\texttt{a},\texttt{b},\texttt{a},\texttt{b}$
Step 1: $\texttt{b,a}\!\to\!\texttt{ba}\,\,\Longrightarrow\,\,\mathbf{v}_{1}=\texttt{ba},\texttt{ba},\texttt{b}$
Step 2: $\texttt{ba,b}\!\to\!\texttt{bab}\,\,\Longrightarrow\,\,\mathbf{v}_{2}=\texttt{ba},\texttt{bab}$
@@ 3.3 Algorithmic Variants
Model developers may want to include and tokenize data for which parallel data is not available or where this concept does not even apply, such as programming code. Also, they may not want to guarantee full parity, but still give a high weight to global compression. We support these goals with a hybrid learning algorithm that uses the global objective of Classical BPE (eq.˜5) for the first $K$ merges, then switches to the parity-aware objective (eq.˜6) for another $J$ merges. $K$ and $J$ can be chosen by model developers to trade off global compression and fairness according to their priorities.
@@ 4.2.1 Intrinsic Metrics
  • Tokenizer Fairness Gini (Meister, 2025) adapts the Gini coefficient to the per-language tokenization cost distribution (e.g., tokens per line (document) in a parallel corpus). Values near 0 mean equal cost across languages; values closer to 1 indicate inequality.
@@
For completeness, we also track Type–Token Ratio and Average Token Rank (vocabulary diversity; Limisiewicz et al., 2023) as well as Rényi entropies (distributional concentration; Zouhar et al., 2023a). We evaluate these metrics using the TokEval suite (Meister, 2025).
@@ 5.1 Intrinsic Evaluation
Table 2: Average downstream performance (accuracy %) across 13 multilingual benchmarks (tokenizers trained on the (unbalanced) 30-lang dataset with $128k$ vocab size). The Random column shows the expected accuracy of a random classifier. Best performance per language is bolded. Benchmark details for each language are provided in ˜E and 7, respectively.
@@ 7 Discussion and Conclusion
Overall, the trade-offs required for using parity-aware BPE are minimal. Inference remains exactly the same as in Classical BPE. During the learning stage, the algorithm adds only an $\mathcal{O}(\mathcal{L})$ pass per merge for recomputing language-level compression rates on a dev corpus, leaving the asymptotic complexity identical to Classical BPE. Further, we observe empirically that a small, sentence-aligned development set is sufficient to drive the “fair-max” decision. When resource or domain mismatches make full equality undesirable, hybrid and moving-window variants further let practitioners trade off global compression versus strict parity; our empirical results validate that these variants perform well in practice. From the model-developer’s perspective, parity-aware BPE is a drop-in replacement: it requires no architectural changes and minimal changes to the tokenizer pipeline.
@@ Acknowledgements
RS acknowledges support by the Swiss National Science Foundation through the MUTAMUR project (no. 213976). We also gratefully acknowledge the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, a Meta LLM Evaluation Research Grant, and the Swiss AI Initiative through a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a06 on Alps. We thank Tiago Pimentel for his feedback on earlier drafts of this manuscript.
@@ B.3 Morphological and Multilingual Fairness Metrics
We use an adaptation of the Gini coefficient—often used as a measure of economic inequality—to encapsulate tokenizer fairness across languages (Meister, 2025). Formally, let $c_{1}\leq c_{2}\leq\ldots\leq c_{n}$ be the “costs” under a given tokenizer $T$ for languages $\mathcal{L}=\{l_{1},l_{2},\ldots,l_{n}\}$. Here, we quantify cost as the average number of tokens it takes to encode the unit of interest (e.g., a byte, word or line); [^9] This is equivalent to fertility, or the inverse of the compression rate. when using a parallel corpus, this can be cost per line (document), which controls for discrepancies between average character byte lengths across different scripts. The Gini coefficient for tokenizer $T$ is then:
@@
Values range from 0 (completely equal costs across languages) to 1 (maximum inequality). This metric condenses multilingual tokenizer fairness into a single number by measuring the degree of inequality in computational costs across languages; lower Gini coefficients indicate more equitable tokenizer compression across languages, while higher values suggest systematic bias toward certain languages.